# INTERACT: Enabling Interactive, Question-Driven Learning in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) excel at answering questions but remain passive learners—absorbing static data without the ability to question and refine knowledge. This paper explores how LLMs can transition to interactive, question-driven learning through student-teacher dialogues. We introduce INTERACT (INTEReractive Learning for Adaptive Concept Transfer), a framework in which a “student” LLM engages a “teacher” LLM through iterative inquiries to acquire knowledge across 1,347 contexts, including song lyrics, news articles, movie plots, academic papers, and images. Our experiments show that across a wide range of scenarios and LLM architectures, interactive learning consistently enhances performance, achieving up to a 25% improvement, with ‘cold-start’ student models matching static learning baselines in as few as five dialogue turns. Interactive setups can also mitigate the disadvantages of weaker teachers, showcasing the robustness of question-driven learning.
@@
Aum Kendapadithanks:   Equal contribution  Kerem Zaman11footnotemark: 1  Rakesh R. Menon11footnotemark: 1  Shashank Srivastava UNC Chapel Hill aum@unc.edu, {kzaman, rrmenon, ssrivastava}@cs.unc.edu
@@ 1 Introduction
  • Analyses showing the importance of adaptive questioning and uncovering insights to improve interactive learning strategies. [^2] Data and code will be shared upon publication.
@@ 2 Related Work
Early work on language-guided machine learning often used single-turn instructions and limited examples (Srivastava et al., 2017; Hancock et al., 2018; Arabshahi et al., 2020; R. Menon et al., 2022). To improve comprehension, researchers have explored active learning (Collins et al., 2008; Tamkin et al., 2022) and language-based clarifications (Rao and Daumé III, 2018; Srivastava et al., 2019). Our approach follows the latter, but unlike methods that rely on templated question generation, we let LLMs produce contextually relevant queries. Our teacher-student paradigm is consonant with knowledge distillation (Hinton, 2015), though here the “student” acquires knowledge by active questioning rather than passively receiving information.
@@ 3.2 Datasets
Our evaluation dataset comprises of 1,347 contexts spanning multiple domains. This compilation includes song lyrics, news articles, movie plots, academic papers and images. Table 1 provides an overview of the domains and the number of concepts per domain in the dataset. Further details about the dataset composition can be seen in Appendix B.
@@
To measure learning performance, we generate a nine-question quiz per concept, with three levels of difficulty and three questions per level. Questions were crafted using gpt-4o-2024-08-06. We used an adversarial filtering strategy to exclude questions that could be answered by a gpt-4o-mini model without reference to the provided context. This ensured that each question required eliciting concept information from the provided context. A manual analysis of a random subset of questions by the authors showed that 97% of them satisfied three criteria: (a) good for testing student understanding, (b) answerable using the context, and (c) do not require deeper knowledge beyond the context. Appendix B includes details of the quiz generation process.
@@ 4.2 RQ2: How Well Can Students Learn Concepts through Interactions?
Figure 4 shows that students without a lesson initially lag behind their static-lesson counterparts, but improve substantially after each question-answer exchange for gpt-4o-mini. These improvements hold across different language models (+22% absolute gain for LLaMA-8B, Table 2; Figure 6 in Appendix C). Despite steady improvement, most dynamic students do not surpass the static-lesson baseline within five rounds, particularly in text domains. Longer or more strategically guided interactions might close the gap further. In the Images domain, however, performance quickly saturates over the first couple of interaction rounds (see Figure 7 for Pixtral-12B and LLaMA-3.2-11B results).
@@ 4.4 RQ4: Can Borrowed Interactions Substitute for Proactive Engagement?
We generate transcripts from an interaction between a strong teacher-student pair (both LLaMA-70B models) and provide them as context (similar to the teacher lesson) to a weaker student (LLaMA-8B) that never engaged in that dialog. We then measure if the weaker student’s performance improves from this passive exposure alone.
@@ 4.5 RQ5: What Interaction Factors Predict Student Learning Gains?
We first collect a comprehensive set of 45 interaction-related features [^6] We provide the full list of features in Appendix Table 9. that can be computed from the teacher-student transcripts. These factors include syntactic features of the student/teacher responses, statistics about tokens, metrics of linguistic complexity, semantic relatedness of questions and responses, among others. For each domain and LM configuration, we extract these features from the recorded interactions and aggregate them into a feature matrix. We then train a random forest regressor, using best-found hyperparameters from cross-validation, to predict the learning gain of the student model after each round of interaction with a teacher. Performance is evaluated using the $R^{2}$ score on a held-out test set.
@@
In most domains, the predictive power of our feature set is low: $R^{2}$ scores on held-out data are often close to zero. However, for the movie plots domain, $R^{2}$ scores reach up to 0.17, suggesting that meaningful signals can be extracted in specific conditions. The top contributing features are: the Flesch reading score of teacher responses, the cumulative token length of all teacher responses up to that point, the length of the student’s questions, and the estimated probability of a teacher response being relevant to the student’s query. This suggests that our current feature sets and metrics are insufficient for robustly capturing the nuanced factors driving interactive learning success. Identifying richer metrics remains a key challenge.
@@ Appendix A Extended Related Work
Appendix B Dataset Creation
B.1 Data Collection
Domain Name
Context
Count
Avg. #
Tokens
Images 150 -
Movie Plots 214 671.3
Song Lyrics 467 296.9
Academic Papers 170 1560.8
   Computer Science 23 1438.5
   Economics 16 1424.2
   Electrical Engineering 25 1443.3
   Mathematics 23 1667.6
   Physics 22 1781.7
   Quantitative Biology 22 1752.1
   Quantitative Finance 15 1462.4
   Statistics 24 1507.1
News Articles 346 1163.7
   Business 38 1068.5
   Entertainment 30 627.7
   Health 28 1372.8
   Politics 43 1690.6
   Science 27 1406.4
   Sports 41 1232.1
   Style 51 1023.0
   Travel 10 1026.6
   US News 28 1166.7
   World News 50 981.7
Total 1,347 967.1
Table 9: Feature Names, Descriptions, and Details
Feature Name Description Details
Question-Level Features
Teacher-Response-Level Features
Interaction-Dynamics Features
Unanswered Queries Count of previously unanswered questions carried forward. Placeholder: Requires QA alignment logic.
Linguistic/Style Features
Semantic/NLP Features
Performance/Contextual Features
Movie Plots
Song Lyrics
News Articles
Academic Papers
Images
B.2 Quiz Questions
Quiz Question Complexities.
Domain Difficulty Example Quiz Questions
Images What color are the flowers in the right garden bed?
College Why does Gary refuse money from Madison initially?
Adversarial Quiz Question Generation.
Quiz Question Validation.
Category # Evaluated Yes (%)
Question 1: Suitable for testing student understanding?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 58 (96.67%)
Song Lyrics 51 48 (94.12%)
Overall 234 227 (97.01%)
Question 2: Answerable using provided context?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 59 (98.33%)
Song Lyrics 51 48 (94.12%)
Overall 234 228 (97.44%)
Question 3: Avoids requiring extra knowledge?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 59 (98.33%)
Song Lyrics 51 48 (94.12%)
Overall 234 228 (97.44%)
Appendix C Additional Results
@@
C.1 Additional Static Results
@@
C.2 Additional Dynamic Results

### v2

@@ Abstract
Large language models (LLMs) excel at answering questions but remain passive learners—absorbing static data without the ability to question and refine knowledge. This paper explores how LLMs can transition to interactive, question-driven learning through student-teacher dialogues. We introduce INTERACT (INTERactive learning for Adaptive Concept Transfer), a framework in which a “student” LLM engages a “teacher” LLM through iterative inquiries to acquire knowledge across 1,347 contexts, including song lyrics, news articles, movie plots, academic papers, and images. Our experiments show that across a wide range of scenarios and LLM architectures, interactive learning consistently enhances performance, achieving up to a 25% improvement, with ‘cold-start’ student models matching static learning baselines in as few as five dialogue turns. Interactive setups can also mitigate the disadvantages of weaker teachers, showcasing the robustness of question-driven learning.  [^1] Our code and dataset are available at https://github.com/aumken/interact.
@@
Aum Kendapadithanks:   Equal contribution  Kerem Zaman11footnotemark: 1  Rakesh R. Menon11footnotemark: 1  Shashank Srivastava UNC Chapel Hill aumken@alumni.unc.edu, {kzaman, rrmenon, ssrivastava}@cs.unc.edu
@@ 1 Introduction
  • Analyses showing the importance of adaptive questioning and uncovering insights to improve interactive learning strategies.
@@ 2 Related Work
Early work on language-guided machine learning often used single-turn instructions and limited examples (Srivastava et al., 2017; Hancock et al., 2018; Arabshahi et al., 2020; R. Menon et al., 2022). To improve comprehension, researchers have explored active learning (Collins et al., 2008; Tamkin et al., 2022) and language-based clarifications (Rao and Daumé III, 2018; Srivastava et al., 2019). Our approach follows the latter, but unlike methods that rely on templated question generation, we let LLMs produce contextually relevant queries. Our teacher-student paradigm is consonant with knowledge distillation (Hinton, 2015), though here the “student” acquires knowledge by active questioning rather than passively receiving information. Our approach is also related to work on document-grounded dialogue Feng et al. (2020, 2021), which model multi-turn interactions grounded in text.
@@
Interactive Learning in Human Tutoring.
Evidence from human learning shows that interactive settings consistently outperform passive instruction. Adaptive tutoring through feedback and dialogue yields substantial gains over classroom-based teaching (Bloom, 1984). These improvements stem not only from personalization, but from the structure of interaction itself, specifically, student-led questioning and responsive explanation (VanLehn, 2011). Our setup parallels this structure: the student model initiates queries to target its own uncertainties. Learning emerges through focused, multi-turn exchanges rather than static instruction.
@@ 3.2 Datasets
Our evaluation dataset comprises of 1,347 contexts spanning multiple domains. This compilation includes song lyrics, news articles, movie plots, academic papers and images. Table 1 provides an overview of the domains and the number of concepts per domain in the dataset. Further details about the dataset composition can be seen in Appendix C.
@@
To measure learning performance, we generate a nine-question quiz per concept, with three levels of difficulty and three questions per level. Questions were crafted using gpt-4o-2024-08-06. We used an adversarial filtering strategy to exclude questions that could be answered by a gpt-4o-mini model without reference to the provided context. This ensured that each question required eliciting concept information from the provided context. A manual analysis of a random subset of questions by the authors showed that 97% of them satisfied three criteria: (a) good for testing student understanding, (b) answerable using the context, and (c) do not require deeper knowledge beyond the context. Appendix C includes details of the quiz generation process.
@@ 4.2 RQ2: How Well Can Students Learn Concepts through Interactions?
Figure 4 shows that students without a lesson initially lag behind their static-lesson counterparts, but improve substantially after each question-answer exchange for gpt-4o-mini. These improvements hold across different language models (+22% absolute gain for LLaMA-8B, Table 2; Figure 6 in Appendix B). Despite steady improvement, most dynamic students do not surpass the static-lesson baseline within five rounds, particularly in text domains. Longer or more strategically guided interactions might close the gap further. In the Images domain, however, performance quickly saturates over the first couple of interaction rounds (see Figure 7 for Pixtral-12B and LLaMA-3.2-11B results).
@@ 4.4 RQ4: Can Borrowed Interactions Substitute for Proactive Engagement?
We generate transcripts from an interaction between a strong teacher-student pair (both LLaMA-70B models) and provide them as context (similar to the teacher lesson) to a weaker student (LLaMA-8B) that never engaged in that particular dialog. We then measure if the weaker student’s performance improves from this passive exposure alone.
@@ 4.5 RQ5: What Interaction Factors Predict Student Learning Gains?
We first collect a comprehensive set of 53 interaction-related features [^6] We provide the full list of features in Appendix Table 12. that can be computed from the teacher-student transcripts. These factors include syntactic features of the student/teacher responses, statistics about tokens, metrics of linguistic complexity, semantic relatedness of questions and responses, among others. For each domain and configuration with gpt-4o-mini as student and teacher models, we extract these features from the recorded interactions and aggregate them into a feature matrix. We then train a random forest regressor, using best-found hyperparameters from cross-validation, to predict the learning gain of the student model after each round of interaction with a teacher. Performance is evaluated using the $R^{2}$ score on a held-out test set.
@@
In most domains, the predictive power of our feature set is low: $R^{2}$ scores on held-out data are often close to zero. However, for the song lyrics domain, $R^{2}$ scores reach up to 0.14, suggesting that meaningful signals can be extracted in specific conditions. The top contributing features are cumulative exposure (number of unique tokens), overlap between quiz questions and student questions, semantic alignment between student and quiz questions, response information density, and response correctness. This suggests that our current feature sets and metrics are insufficient for robustly capturing the nuanced factors driving interactive learning success. Identifying richer metrics remains a key challenge.
@@ 5 A Future of Conversational Learning
6 Acknowledgements
This work was supported by the NSF grant DRL2112635. The authors would also like to thank anonymous reviewers for valuable feedback to improve the manuscript.
@@ Appendix A Extended Related Work
Appendix B Additional Results
B.1 Interaction Utilization
As shown in Section 4.4, passive exposure to high-quality content is insufficient to replace proactive engagement. Specifically, providing transcripts of interactions between a stronger teacher-student pair to a weaker student does not lead to comparable performance. We also observe that interactions involving weaker students do not substantially degrade overall performance. To assess whether models genuinely leverage the observed interactions, we conduct an ablation study in which we replace meaningful interactions with irrelevant ones.
Student $\rightarrow$ w/o Lesson w/ Lesson
Interactions $\rightarrow$ Orig. Random Orig. Random
Academic Papers 78.54 47.33 (-31.21) 82.56 47.17 (-35.39)
Movie Plots 57.85 34.11 (-23.74) 73.94 34.21 (-39.73)
News Articles 87.40 73.11 (-14.29) 92.52 72.95 (-19.57)
Song Lyrics 68.92 45.84 (-23.07) 74.94 45.53 (-29.41)
Images 49.47 23.73 (-25.73) 87.07 23.60 (-63.47)
Table 8: Quiz performance comparison for gpt-4o-mini with its original and random interactions for students without/with lessons. Random interaction results are averaged across 3 seeds.
Study Design.
For each domain, we sample 100 concept interactions and replace them with interactions from a randomly selected, unrelated concept within the same domain. In this setup, the interactions are sourced from dialogues between gpt-4o-mini student and teacher models.
Main Results.
Table 8 shows that exposure to irrelevant transcripts results in a significant performance decline. This indicates that students are indeed making effective use of the observed interactions.
@@ B.1 Interaction Utilization
B.2 Additional Static Results
@@
B.3 Additional Dynamic Results
@@
Appendix C Dataset Creation
C.1 Data Collection
Domain Name
Context
Count
Avg. #
Tokens
Images 150 -
Movie Plots 214 671.3
Song Lyrics 467 296.9
Academic Papers 170 1560.8
   Computer Science 23 1438.5
   Economics 16 1424.2
   Electrical Engineering 25 1443.3
   Mathematics 23 1667.6
   Physics 22 1781.7
   Quantitative Biology 22 1752.1
   Quantitative Finance 15 1462.4
   Statistics 24 1507.1
News Articles 346 1163.7
   Business 38 1068.5
   Entertainment 30 627.7
   Health 28 1372.8
   Politics 43 1690.6
   Science 27 1406.4
   Sports 41 1232.1
   Style 51 1023.0
   Travel 10 1026.6
   US News 28 1166.7
   World News 50 981.7
Total 1,347 967.1
Table 12: Feature Names, Descriptions, and Details
Feature Name Description Details
Question-Level Features
Bloom’s Taxonomy Corresponding Bloom’s taxonomy level (Bloom and Krathwohl, 1966). A binary feature for each level. Utilizes gpt-4o to decide the taxonomy.
Teacher-Response-Level Features
Response Sentiment Sentiment score for the response. Utilizes Hugging Face’s text classification pipeline (clapAI/roberta-base-multilingual-sentiment)
Interaction-Dynamics Features
Unanswered Queries Count of unanswered questions. Uses gpt-4o to determine whether a given response answers the corresponding question.
Student Context Coverage The cumulative overlap between question entities so far and context entities. Utilizes spaCy’s NER to extract and count unique entities.
Teacher Context Coverage The cumulative overlap between response entities so far and context entities. Utilizes spaCy’s NER to extract and count unique entities.
Student Quiz Coverage The cumulative overlap between question tokens so far and quiz question tokens. The ratio of unique quiz tokens covered by the questions.
Teacher Quiz Coverage The cumulative overlap between response tokens so far and quiz question tokens. The ratio of unique quiz tokens covered by the responses.
Student Semantic Alignment The average maximum similarity between each quiz question and the questions asked so far. Cosine similarity of embeddings.
Teacher Semantic Alignment The average maximum similarity between each quiz question and the responses so far. Cosine similarity of embeddings.
Linguistic/Style Features
Semantic/NLP Features
Performance/Contextual Features
Instruction-Following Score Adherence to teacher instructions. Uses reward model scores as proxy. (Ray2333/GRM-Gemma2-2B-rewardmodel-ft)
Movie Plots
Song Lyrics
News Articles
Academic Papers
Images
C.2 Quiz Questions
Quiz Question Complexities.
Domain Difficulty Example Quiz Questions
Images What color are the flowers in the right garden bed?
College Why does Gary refuse money from Madison initially?
Adversarial Quiz Question Generation.
Quiz Question Validation.
Category # Evaluated Yes (%)
Question 1: Suitable for testing student understanding?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 58 (96.67%)
Song Lyrics 51 48 (94.12%)
Overall 234 227 (97.01%)
Question 2: Answerable using provided context?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 59 (98.33%)
Song Lyrics 51 48 (94.12%)
Overall 234 228 (97.44%)
Question 3: Avoids requiring extra knowledge?
Academic Papers 72 70 (97.22%)
Movie Plots 51 51 (100.0%)
News Articles 60 59 (98.33%)
Song Lyrics 51 48 (94.12%)
Overall 234 228 (97.44%)